CHIRPED-PULSED FTMW SPECTRUM OF VALERIC ACID AND 5-AMINOVALERIC ACID. A STUDY OF AMINO ACID MIMICS IN THE GAS PHASE^{a}

\underline{RYAN G. BIRD}, VANESA VAQUERO, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, Pa 15213; JUSTIN L. NEILL and BROOKS H. PATE, Department of Chemistry, University of Virginia, Charlottesville, Va 22904.

Microwave studies of the structural and dynamical properties of several organic acids and their water complexes have been described by a number of research groups. Here we continue this theme by the study of valeric acid and 5-aminovaleric acid, using chirped-pulsed Fourier transform microwave spectroscopy (CP-FTMW). The rotational spectrum from 6.5 to 18 GHz was collected using a compilation of 250 MHz chirped pulses and pieced together. Their structures and water complexes were determined and will be compared to other amino acids.

aWork supported by NSF (CHE-0618740 and -0911117).